### v1

@@ 5.2 Prompt Analysis
This two-step process allows the model to provide an explanation for its decision and leverage the reasoning and examples to produce a more accurate and robust output. This superiority can be quantified: for example, in the textual modality, the ACoT base achieves a 51.40% accuracy, which is approximately 81% higher than the ZS base (28.40%) and 8.4% higher than the FS base (47.40%). In terms of the F1 score, the ACoT base (50.40%) outperforms the ZS base (42.40%) by 18.9% and the FS base (48.40%) by about 4.1%. This difference is less pronounced in numerical tasks, where the ACoT base achieves an accuracy of 40.40%, 47.4% higher than the ZS base (27.40%), and 8% higher than the FS base (37.40%).
@@
These differences can be attributed to how these methods prompt the model. ZS and FS methods direct the model to generate answers straight from the input without guiding it through intermediate reasoning. While this may suffice for more straightforward tasks, it often falls short when dealing with more complex tasks that require deeper understanding.
@@
ACoT, by contrast, excels in these scenarios because it prompts the model to break down the task into a series of logical, intermediate steps before concluding. This approach mirrors the “self-explanation” cognitive strategy, where breaking down and explaining each step improves problem-solving abilities. This explains why ACoT significantly outperforms ZS and FS, especially in tasks that demand nuanced understanding and intermediate reasoning.
@@
Prompt Modality Mode Accuracy Precision Recall F1
ZS Textual Base 28.40 27.30 98.30 42.40
Textual Finetuned 28.40 27.30 99.30 42.40
Numerical Base 27.40 27.10 98.70 43.60
Numerical Finetuned 27.20 27.80 98.40 42.60
FS Textual Base 47.40 33.80 91.30 48.40
Textual Finetuned 46.40 30.80 97.30 45.40
Numerical Base 43.40 31.80 92.30 46.20
Numerical Finetuned 45.40 32.10 87.30 46.60
ACoT Textual Base 51.40 34.80 90.30 50.40
Textual Finetuned 46.40 31.80 85.30 47.60
Numerical Base 40.40 31.20 93.30 41.60
Numerical Finetuned 40.10 32.80 84.30 40.60
@@
5.3 Finetuning Analysis
We analyze the performance of LLAMA 3.1 8B when fine-tuned on Assembly101-O (Table 4). Contrary to the common expectation that fine-tuning enhances performance by adapting the model to the task’s specific characteristics and distribution, this is not consistently observed. In both the ZS and FS methods, fine-tuning offers no significant improvement across metrics, with some cases showing similar or even lower precision and recall values. Notably, in the ACoT method, fine-tuning consistently results in a decrease in performance, particularly in terms of precision, suggesting that fine-tuning may not always provide the anticipated benefits.
@@ 5.4 Speed Analysis
Model Speed (s/sample)
@@
6 Aggregation Strategy
6.1 Step Recognizer Analysis
The action recognizer is the architecture module responsible for online action detection, a task extensively explored in recent research due to its various possible applications. Several models have been proposed using datasets such as Thumos [52], TVSeries [53], and Fine Action [54], with the current leading model being MiniROAD [32]. We rely on the state-of-the-art MiniROAD, even though it should be considered that its performances on procedural datasets has been suboptimal, largely due to two key factors: the shift in action distribution between the training and testing phases, and the inherent class imbalance within the dataset. The distribution shift occurs because the actions seen during training often differ from those encountered during testing, leading to poorer generalization. Additionally, Assembly101 contains a significant imbalance in class frequencies (see Figure 4), where common actions dominate the dataset while less frequent yet crucial procedural actions are underrepresented.
Figure 4: 20 most frequent classes in Assembly101-O
6.2 Frame Aggregation
@@
To address this, we propose different frame aggregation strategies to efficiently convert per-frame predictions into coherent action sequences, aiming to reduce noise and stabilize predictions. This results in a cleaner input for the anticipator module. Our approach introduces three distinct aggregation methods, as shown in Figure 5.
@@
Figure 5: Predictions bars: (a) Ground truth predictions, (b) Predictions using the recognizer,(c) First aggregation strategy, (d) Second aggregation strategy, (e) Third aggregation strategy
@@
Each aggregation strategy’s performance will be evaluated using the Levenshtein Similarity between the predicted and ground truth action sequence, which derives from the Levenshtein distance. The similarity between two action sequences A and B is defined as:
Levenshtein $\text{Levenshtein Similarity}(A,B)=1-\frac{d_{dev}(A,B)}{\max(len(A),len(B))}$ (4)
@@
where: $d_{dev}(A,B)$ is the Levenshtein distance between procedure A and B, which is the minimum number of single-action insertions, deletions, or substitutions required to transform procedure A into procedure B, and len(A) and len(B) represent the length of procedure A and B, respectively. Note that the term max(len(A),len(B)) normalizes the score by the length of the longer procedure, ensuring that the ratio is between 0 and 1.
6.2.1 First strategy
@@
The first aggregation strategy (Figure 5.c) we propose consists of a two-step process. In the first step, we calculate the mode for each non-overlapping sliding window of frames of a given length $l$. Because the windows are non-overlapping, this approach introduces a slight delay in processing, which is a trade-off for improved prediction stability. The mode, representing the most frequent action label within that window, replaces all predictions in that window, ensuring consistency across the frames and removing the noise of smaller portions wrongly predicted. In the second step, we apply the elimination of successive duplicates to remove consecutive repeated action labels, which prevents redundant predictions. This cleaned output serves as the input to the anticipation module, ultimately aiming to produce a more stable and accurate sequence of action predictions.
6.2.2 Second strategy
@@
The second aggregation strategy (Fig. 5.d) we propose employs a sliding window approach with a stride of $1$ frame, allowing for overlapping portions of frames. In this method, we evaluate the mode within the overlapping window that includes the last frame and the previous $l-1$ frames. The prediction of the last frame is then substituted with this mode, which captures the most frequent action label in that localized area. This method is designed with the intent to reduce delay, as it processes frames more dynamically compared to non-overlapping windows.
6.2.3 Third Strategy
@@
The third aggregation strategy (Fig. 5.e) uses a stride of 1 frame, similar to the previous approach. To address the potential issue of the mode being incorrectly positioned within the window, we opt to substitute the prediction of the central frame rather than the final frame, with some corrections for the initial and final frames in the video. This adjustment ensures that the most frequent action label is applied to a frame that is more representative of the overall context, thus enhancing the accuracy of the prediction.
@@
Figure 6: Levenshtein similarity for the proposed aggregation strategies. The maximum similarity is achieved by the first strategy when using a window size of 500 frames.
Figure 7: Histogram of the durations of the action in Assembly101-O
Figure 6 illustrates that the first approach is the most effective in maximizing similarity.

### v2

@@ 3.3 Step Anticipation
{contextframe}
Contextual Prompt:
Toy: “bulldozer”
Input Sequence:
attach-cabin, attach-body, attach-track, attach-blade
Next Symbol:
attach-figurine
{sequenceframe}
Current sequence:
attach-door
3.4 Mistake Detection
Finally, we analyze the output of the two modules to identify procedural errors. Specifically, we classify steps as correct when the outputs of both modules match, while we label cases as errors when the outputs differ. That is:
$\begin{cases}s^{\rho}_{\tau}\neq s^{\xi}_{\tau}&\text{MISTAKE}\\ s^{\rho}_{\tau}=s^{\xi}_{\tau}&\text{CORRECT}\\ \end{cases}$ (2)
4 Experiments
This section describes the benchmark datasets, adopted evaluation metrics and the main experimental results. We introduce Assembly101-O and Epic-tent-O as adaptations of the original datasets [2, 1], detailing the selected labeling criteria for online benchmarking, a novel configuration of training and test splits to accommodate open-set procedural mistakes. We define our proposed online metrics in Sec. 4.2. Finally, we introduce the main experimental results in Sec. 4.3, whereas further insights that motivate our design choices are provided through the ablation studies in Sec. 5.
4.1 Datasets
To align with real-world applications, this work focuses on Egocentric videos.
4.1.1 Assembly101-O
Assembly101 [2] is a large-scale video dataset containing 362 videos of assembly and disassembly tasks involving 101 types of toy vehicles, recorded from eight multiple static and four egocentric cameras. It includes extensive annotations at various levels of granularity, with over 100K coarse action segments and 1M fine-grained action segments and 18M 3D hand poses. The dataset supports diverse challenges, including action anticipation and segmentation, mistake detection and 3D pose-based action recognition.
Assembly101 for Online and Open-Set Mistake Detection (Proposed) We introduce a novel split of the dataset [2] tailored explicitly for online, open-set mistake detection. Assembly101-O introduces two main modifications to [2] as we show in Figure 5: a new train/test split and revised procedure lengths.
The new split assigns all correct procedures to the training set while reserving videos with mistakes for testing. This adjustment enables models to learn sequences characteristic of correct procedures in a one-class classification (OCC) framework, avoiding bias from specific mistake types during training.
Figure 5: The modification occurs in two steps: starting from Assembly101 (a), first, a new train/test split assigns all correct procedures to the training set, reserving videos with mistakes for testing (b). Second, the lengths of procedures containing mistakes are adjusted by trimming them to the first mistake, preventing the creation of corrupted sequences (c). This setup enables models to learn correct sequences within a one-class classification (OCC) framework, treating any deviation as a mistake and ensuring a balanced test set for effective mistake detection.
The second adjustment limits each procedure’s evaluation to the point where a mistake caused by an incorrect step that distrupts the workflow: evaluating beyond the point of procedural compromise creates discrepancies between the training and test set procedures, hindering effective step recognition and anticipation.
4.1.2 Epic-tent-O
Epic-tent is a dataset of egocentric videos capturing the outdoor assembly of a camping tent. It was collected by 24 participants who wore two head-mounted cameras (a GoPro and a SMI eye tracker) while performing the task, resulting in 5.4 hours of recorded video. The dataset includes annotations for action labels, task errors, participant self-rated uncertainty and gaze positions, capturing the variability and complexity of the task as participants interacted with non-rigid objects such as the tent, guylines, instructions and tent bag, and displayed varying proficiency and confidence levels.
Epic-tent for Online and Open-Set Mistake Detection (Proposed) We propose a novel split for the Epic-tent dataset [1] tailored for open-set mistake detection. The dataset includes annotations for nine distinct mistake types, though only four of these categories —“order”, “omit”, “correction” and “repeat” — are considered procedural mistakes, as they represent deviations that compromise the task’s procedural integrity. Other categories such as “slow”, “search”, “misuse”, “motor” and “failure” are not procedural errors, as they do not disrupt the task’s progression.
Unlike Assembly101 [2], Epic-tent is designed explicitly for supervised error detection with each recorded procedure containing at least one mistake, making it incompatible with the split procedure proposed for Assembly101-O. However, Epic-tent provides confidence scores from each participant, reflecting their self-assessed uncertainty during task execution. Therefore, we define a custom split strategy where videos from higher-confidence participants form the training set. In contrast, participants showing more significant uncertainty, and thus potentially prone to more errors, make up the test set. In practice, the videos that form the test set have a median confidence score under 0.6, while the others form the train set. Only 22 videos have the confidence score annotations, while the remaining 7 do not and are assigned to the test set. This partition results in 14 training videos and 15 test videos.
This strategy is particularly promising for real-world applications where accurately labeling erroneous frames may be challenging or training a mistake detector can start soon after recording without requiring full annotation completion.
Videos in the test set are trimmed to the frame containing the first procedural mistake as in Assembly101-O, while videos representing correct procedures remain unaltered.
4.2 Evaluation Definition
To assess the performance of our procedural mistake detection model, we consider True Positives (TP) as the count of errors correctly identified, and True Negatives (TN) as the count of correctly performed steps that are not flagged as errors. Since, in ego-centric mistake detection, the positive class (actual mistakes) is typically much smaller than the negative class (correct steps), a straightforward application of classical Precision will tend to underestimate the impact of False Positives (FP). In particular, even a small FP rate can yield a large absolute number of FP simply because there are so many more negative (non-error) samples than positive (error) samples.
To mitigate this imbalance, we utilize Balanced Precision [23], denoted $\mathrm{Prec}_{b}$, by scaling the contribution of each False Positive by the ratio of positive-to-negative samples. Concretely, if $N_{+}$ is the total number of positive (error) samples and $N_{-}$ is the total number of negative (correct) samples, we define:
$\mathrm{Prec}_{b}\;=\;\frac{\mathrm{TP}}{\mathrm{TP}\;+\;\Bigl{(}\tfrac{N_{+}}{N_{-}}\Bigr{)}\,\mathrm{FP}}\,.$ (3)
Thus, FP is down-weighted by the factor $\tfrac{N_{+}}{N_{-}}$, ensuring that the penalty for mistakenly labeling a correct step as an error is placed on the same scale as the reward for correctly identifying a mistake. When $N_{+}\ll N_{-}$, this adjustment prevents a handful of FPs from dominating the denominator and driving classical Precision toward zero.
We retain Recall (also known as Sensitivity for the positive class) in its standard form, since Recall inherently measures the proportion of actual mistakes that the model retrieves:
$\mathrm{Recall}\;=\;\frac{\mathrm{TP}}{\mathrm{TP}+\mathrm{FN}}\,.$ (4)
Because Recall does not depend on the cardinality of the negative class, it is not directly skewed by the negative-class majority and therefore needs no further adjustment.
[…]
4.3 Experiments
4.3.1 Compared Methods
To estimate the effectiveness of our model, we evaluate its performance by comparing it against several baseline models based on the metrics presented in Sec. 4.2:
One-step memory The One-step memory baseline stems from a transition matrix that considers only the correct procedures. Specifically, given the set of actions $\mathcal{A}$ in the training set with $|\mathcal{A}|=C$, we define a transition matrix $M\in\mathbb{R}^{C\times C}$, which records, in position $(l,m)$, the number of occurrences where action $m$ follows action $l$. During evaluation, an action in the test split that does not correspond to a transition recorded in this training set matrix is labeled as a mistake. This baseline provides a simple method for recognizing deviations from standard procedural transitions.
BERT-based Mistake Detection We also leverage the capabilities of BERT [56], using its specific [CLS] token to predict whether an action sequence is correct or erroneous. More specifically, we fine-tune BERT with a next-token-prediction task, training the model to determine whether step B logically follows step A within a given procedure. In our setting, each step is represented as a set of two words, such as attach wheel, which describes coarse-level actions. During evaluation, BERT is presented with pairs of actions and predicts whether the sequence of those actions is correct. The rationale of BERT lies in its pre-training on a vast corpus of text, which helps it understand contextual relationships between procedural steps.
Learned Task Graph Representations We also evaluate an approach based on explicit task graph representations [16]. Procedural activities consist of sequences of key steps that work toward specific goals, making task graphs a valuable, human-understandable representation for capturing the partial ordering of these steps. Unlike traditional methods that rely on hand-crafted procedures to extract task graphs from videos, this approach uses maximum likelihood optimization of edge weights to directly learn task graphs, allowing gradient-based learning that can integrate seamlessly with neural networks. While effective, this baseline is not directly comparable to our method, as it requires access to the entire dataset during training to extract the task graph, whereas our approach relies only on few-shot inference for anticipating actions. We refer to the “Direct Optimization” method trained with the “Task-Graph Maximul Likelihood” loss proposed in [16] as TGML-DO.
4.3.2 Results
[…]
The TGML-DO method [16], which learns task graph representations through gradient-based optimization, also demonstrates strong performance. However, unlike our anticipator, which operates in a few-shot inference setting, TGML-DO checks whether the observed action is correct by referencing domain-specific procedural knowledge mined from the entire dataset during training. This distinction likely contributes to the superior generalization of our model, particularly when using predicted steps instead of oracle labels in the Assembly101-O dataset. Our method proves to be more robust to procedural variations and better suited for handling previously unseen cases.
[…]
A more detailed discussion on the effectiveness of LLMs for step anticipation is provided in Section 5.1.
Assembly101-O Epic-tent-O
Step Recog. Step Antic. Precisionb Recall F1 scoreb Precisionb Recall F1 scoreb
One-step memory Oracle 43.87 30.7 36.2 42.1 26.6 32.6
BERT [56] Oracle 93.2 20.0 32.9 71.6 5.6 10.4
PREGO Oracle GPT-3.5 51.4 87.5 64.7 25.6 90.6 39.9
PREGO [22] Oracle DeepSeek-R1 51.7 98.3 67.8 23.8 100 38.4
TI-PREGO Oracle Llama 3.1 60.4 97.8 74.7 40.7 100 57.8
TGML-DO [16] Oracle 77.82 90.4 83.64 91.24 27.3 42.03
TGML-DO [16] MiniRoad [37] 53.77 37.3 44.05 96.55 14.1 24.60
OadTR for MD [14] OadTR [14] OadTR [14] 50.9 18.1 26.7 40.4 21.7 28.3
PREGO MiniRoad [37] GPT-3.5 50.2 75.8 61.4 19.5 73.3 29.4
PREGO [22] MiniRoad [37] DeepSeek-R1 51.2 96.5 65.7 20.3 93.3 33.6
TI-PREGO MiniRoad [37] Llama 3.1 51.5 97.2 67.4 21.2 93.3 34.6
5 Ablation
In the following subsections, we investigate the application of large-language models as step anticipators 5.1. This subsection covers our model selection process, a detailed analysis of different prompting methods like Zero-Shot, Few-Shot, and Automatic-Chain-of-Thought (ACoT), and the effects of fine-tuning and speed analysis. Together, these experiments illustrate the strengths and challenges of using LLMs for action anticipation and highlight the trade-offs between accuracy, reasoning complexity and computational efficiency.
We also present a comprehensive evaluation of the performance of our system and the various strategies used to improve action prediction in assembly tasks. We introduce our Aggregation Strategy 5.2, detailing three methods to convert noisy per-frame predictions into coherent action sequences. We discuss the design of each strategy and their trade-offs, such as delay versus stability, and assess their effectiveness using the Levenshtein similarity metric.
5.1 Large Language Models as Step Anticipators
This section comprehensively analyzes how an action anticipation based on large language models (LLMs) operates using various techniques and methodologies. We begin by evaluating the performance of several LLMs using in-context learning with few-shot examples (Sec. 5.1.1), which allows us to determine which model performs best in this setting. Once the most effective LLM is identified, we investigate different prompting strategies, starting from a system prompt’s impact, including zero-shot, few-shot, and Automatic-Chain-of-Though prompting (Sec 5.1.2). Furthemore, we explore the potential improvements by fine-tuning each prompting method to enhance the model’s performance (Sec. 5.1.3). Lastly, we perform a speed analysis, comparing different prompting methods (Sec. 5.1.4).
All experiments use ground truth labels of the Assembly101-O dataset as the output of the predictor’s branches. This approach ensures that our performance analysis remains objective and untainted by errors from the step recognition branch, allowing for a more accurate evaluation of the step anticipator’s capabilities.
5.1.1 Model Selection
[…]
The superior performance of LLAMA 3.1 8B could be attributed to several factors, including potential advancements in its training methodology, the quality and diversity of its training data, and possible optimizations for general task comprehension. A pattern observed across all models is the substantial disparity between $\mathrm{\text{Precision}}_{b}$ and Recall scores. The consistently high Recall scores indicate that these models identify relevant assembly steps effectively, but the corresponding drop in $\mathrm{\text{Precision}}_{b}$ suggests they also introduce a non-negligible number of spurious steps, underscoring the need for strategies, such as refined prompt engineering.
Table 3: Results of different LLMs on Assembly101-O.
Model Precisionb Recall F1b
Mistral 7B 53.4 88.9 68.1
Gemma 9B 51.5 85.6 65.5
Phi 3 medium 4k instruct 54.0 88.7 68.4
LLAMA 2 7B 54.0 94.0 68.6
LLAMA 3 7B 53.8 87.3 67.3
DeepSeek-R1 51.5 97.8 67.5
LLAMA 3.1 8B 56.7 91.3 69.4
[…]
5.1.2 Prompt Analysis
System prompt.
The results in Table 4 highlight the positive impact of combining Automatic-Chain-of-Though (ACoT) reasoning with Few-Shot (FS) examples on the performance of LLAMA 3.1 8B across different modalities in the Assembly101-O dataset. In all prompting scenarios, the system prompt is written as: {systemframe} System: I am going to provide an input sequence that represents a sequence of actions. Your task is to predict the next action of the last sequence based on the patterns observed in the provided input. Limit yourself to only answer with the predicted sequence, and follow the same format given as input.
This prompt consistently offers essential context about the task, improving the model’s performance, particularly in precision.
In Zero-Shot (ZS) settings, the instruction prompt is presented as: {systemframe} System: Below is an instruction that describes the task, paired with an input. Write a response that appropriately completes the request.
We explore two different input representations: a textual and a numerical one. In the first, we represent an action by its name and natural language, e.g., attach-wheel; in the second one, we use the index associated with the action triplet.
{contextframe}
Textual Representation:
Toy:
“bulldozer”
Input Sequence:
attach-cabin, attach-body, attach-track, attach-blade
Next Symbol:
attach-figurine
{contextframe}
Numerical Representation:
Toy:
“bulldozer”
Input Sequence:
5678, 91011, 1213, 1415
Next Symbol:
1617
[…]
Table 4: Results of LLAMA 3.1 8B on Assembly101-O adding a system prompt (with corrected Precision and Recall values).
System Prompt Modality Precisionb Recall F1 scoreb
Textual 56.7 89.6 69.4
Numerical 54.5 90.3 67.9
X Textual 54.9 86.5 67.1
Numerical 52.4 94.6 65.7
Prompting Methods.
When comparing ZS, FS, and ACoT prompting methods, ACoT outperforms the rest of the prompting techniques (Table 5), particularly when using the pre-trained model (Base), across both textual and numerical tasks. In the ACoT approach, automatic ACoT reasoning is employed, where the LLM is queried twice: first, to generate an internal reasoning step for the answer, and then to use this reasoning, combined with a few examples (as in FS), to derive the final answer. The first query is:
{acotframe}
Let’s analyze each action in detail:
  1. 1.
    For the action “attach the cabin” consider the immediate effect and long-term consequences.
  2. 2.
    Repeat this analysis for each action in the sequence.
  3. 3.
    Identify patterns and causal relationships between the actions.
Now, let’s proceed with the analysis step-by-step:
@@ 5.1.2 Prompt Analysis
This two-step process allows the model to provide an explanation for its decision and leverage the reasoning and examples to produce a more accurate and robust output. In terms of the F1 score, the ACoT base (74.7%) outperforms the ZS base (61.30%) by 19.7% and the FS base (69.50%) by 7.2%.
@@
These differences can be attributed to how these methods prompt the model. ZS and FS methods direct the model to generate answers straight from the input without guiding it through intermediate reasoning. Although this may suffice for more straightforward tasks, it often falls short when dealing with more complex tasks that require deeper understanding.
@@
ACoT, in contrast, excels in these scenarios because it prompts the model to break down the task into a series of logical intermediate steps before concluding. This approach mirrors the “self-explanation” cognitive strategy, where breaking down and explaining each step improves problem-solving skills. This explains why ACoT significantly outperforms ZS and FS, especially in tasks that demand nuanced understanding and intermediate reasoning.
@@
Prompt Modality Mode Precisionb Recall F1 scoreb
ZS Textual Base 47.1 87.9 61.3
Textual Finetuned 48.4 93.4 63.8
Numerical Base 49.1 93.4 64.4
Numerical Finetuned 48.4 91.2 63.2
FS Textual Base 56.8 89.6 69.5
Textual Finetuned 53.1 94.5 68.0
Numerical Base 54.5 90.1 67.9
Numerical Finetuned 55.4 86.9 67.6
ACoT Textual Base 60.4 97.8 74.7
Textual Finetuned 56.3 87.9 68.6
Numerical Base 50.5 83.0 62.8
Numerical Finetuned 51.0 72.5 59.9
@@
5.1.3 Finetuning Analysis
We analyze the performance of LLAMA 3.1 8B when fine-tuned on Assembly101-O (Table 5). Contrary to the common expectation that fine-tuning enhances performance by adapting the model to the task’s specific characteristics and distribution, this is not consistently observed. In both the ZS and FS methods, fine-tuning offers no significant improvement across metrics, with some cases showing similar or even lower $\mathrm{\text{Precision}}_{b}$ and Recall values. Notably, in the ACoT method, fine-tuning consistently results in a decrease in performance, particularly in terms of  $\mathrm{\text{Precision}}_{b}$, suggesting that fine-tuning may not always provide the anticipated benefits.
@@ 5.1.4 Speed Analysis
Model Speed (sample/s)
@@
5.2 Aggregation Strategy
@@
We evaluated three distinct frame aggregation strategies (Fig. 6) aimed at refining per-frame predictions into consistent action sequences, effectively reducing noise and enhancing stability. This results in a cleaner input for the anticipator module.
@@
Figure 6: Predictions bars: (a) Ground truth predictions, (b) Predictions using the recognizer,(c) NOMA: Non-Overlapping Mode Aggregation, (d) OMA: Overlapping Mode Aggregation, (e) OCMA: Overlapping Centered Mode Aggregation
@@
Each aggregation strategy’s performance was evaluated using the Levenshtein Similarity between the predicted and ground truth action sequence, which derives from the Levenshtein distance.
The similarity between two action sequences (i.e procedures) A and B is defined as:
Levenshtein $\text{Levenshtein Similarity}(A,B)=1-\frac{d_{Lev}(A,B)}{\max(len(A),len(B))}$ (6)
@@
where: $d_{Lev}(A,B)$ is the Levenshtein distance between A and B, which is the minimum number of single-action insertions, deletions or substitutions required to transform procedure A into procedure B, and $len(A)$ and $len(B)$ represent the length of procedure A and B, respectively. Note that the term $max(len(A),len(B))$ normalizes the score by the length of the longer procedure, ensuring that the ratio is between 0 and 1.
Non-Overlapping Mode Aggregation (NOMA)
@@
The first aggregation strategy we propose (Figure 4) consists of a two-step process. In the first step, we calculate the mode for each non-overlapping sliding window of frames of a given length $L$. Because the windows are non-overlapping, this approach introduces a slight delay in processing, which is a trade-off for improved prediction stability. The mode, representing the most frequent action label within that window, replaces all predictions in that window, ensuring consistency across the frames and removing the noise of smaller portions wrongly predicted. In the second step, we apply the elimination of successive duplicates to remove consecutive repeated action labels, which prevents redundant predictions. This cleaned output serves as the input to the anticipation module, ultimately aiming to produce a more stable and accurate sequence of action predictions.
Overlapping Mode Aggregation (OMA)
@@
The second aggregation strategy we propose (Fig. 4) employs a sliding window approach with a stride of $1$ frame, allowing for overlapping portions of frames. In this method, we evaluate the mode within the overlapping window that includes the last frame and the previous $L-1$ frames. The prediction of the last frame is then substituted with this mode, which captures the most frequent action label in that localized area. This method is designed with the intent to reduce delay, as it processes frames more dynamically compared to non-overlapping windows.
Overlapping Centered Mode Aggregation (OCMA)
@@
The third aggregation strategy (Fig. 4) uses a stride of 1 frame, similar to the previous approach. To address the potential issue of the mode being incorrectly positioned within the window, we opt for substituting the prediction of the central frame rather than the final frame, with some corrections for the initial and final frames in the video. This adjustment ensures that the most frequent action label is applied to a frame that is more representative of the overall context, thus enhancing the accuracy of the prediction.
@@
Figure 7: Levenshtein similarity for the proposed aggregation strategies. The maximum similarity is achieved by the first strategy when using a window size of 200 frames.